ProbabilityMCQMTP Jun 24 Series IQuestion 3367 of 187
All Questions

Ram is known to hit a target in 2\displaystyle 2 out of 3\displaystyle 3 shots whereas Shyam is known to hit the same target in 5\displaystyle 5 out of $11$ shots. What is the probability that the target would be hit if they both try.

Select Your Answer (Keyboard: A, B, C, D)
For any discrepancies in this question, email contact@cadada.in

Correct Answer

✅ Option a — 9/11\displaystyle 9/11

All Options:

  • A9/11\displaystyle 9/11
  • B6/11\displaystyle 6/11
  • C$10/33$
  • D1/11\displaystyle 1/11

Detailed Solution & Explanation

**Probability that the Target is Hit** Let R\displaystyle R and S\displaystyle S be the events that Ram and Shyam hit the target, respectively. Given: - P(R)=23⇒P(Rˉ)=13\displaystyle P(R) = 23\displaystyle \frac{2}{3} \Rightarrow P(Rˉ\displaystyle \bar{R}) = 13\displaystyle \frac{1}{3} - P(S)=511⇒P(Sˉ)=611\displaystyle P(S) = 511\displaystyle \frac{5}{11} \Rightarrow P(Sˉ\displaystyle \bar{S}) = 611\displaystyle \frac{6}{11} The target is hit if at least one of them hits it. Using the complement rule: P(hit)=1−P(neither hits)P(hit\displaystyle \text{hit}) = 1 - P(neither hits\displaystyle \text{neither hits}) Assuming their attempts are independent: P(neither hits)=P(Rˉ)×P(Sˉ)=13×611=211P(neither hits\displaystyle \text{neither hits}) = P(Rˉ\displaystyle \bar{R}) ×\displaystyle \times P(Sˉ\displaystyle \bar{S}) = 13\displaystyle \frac{1}{3} ×\displaystyle \times 611=211\displaystyle \frac{6}{11} = \frac{2}{11} Now, compute the probability that the target is hit: P(hit)=1−211=911P(hit\displaystyle \text{hit}) = 1 - 211=911\displaystyle \frac{2}{11} = \frac{9}{11} Hence, **Option A** is the correct answer.

About This Chapter: Probability

Paper

Paper 3: Quantitative Aptitude

Weightage

5-7 Marks

Key Topics

Probability Operations, Expected Value

A logic-heavy chapter dealing with random experiments, events (mutually exclusive, exhaustive), set theory probability, conditional probability, and Bayes' Theorem. It forms the basis for Theoretical Distributions.

View Official ICAI Syllabus

Exam Strategy Tip

Always draw a quick Venn Diagram or tree when faced with 'At least one' or 'Only A but not B' wording. It saves you from double-counting.

Key Concepts to Understand

More Questions from Probability

If P(A)=12\displaystyle P(A) = 12\displaystyle \frac{1}{2}, P(B)=13\displaystyle P(B) = 13\displaystyle \frac{1}{3}, P(A∩B)=14\displaystyle P(A \cap B) = 14\displaystyle \frac{1}{4}, then the value of P(A′∪B′)\displaystyle P(A' \cup B') is

Two dice are thrown simultaneously. Find the probability that the sum of digits on the two dice would be 8 or more.

A number is selected from the first 20 natural numbers. Find the probability that it would be divisible by 3 or 7.

A father had three sons namely, Kailash, Harish and Prakash. All are above 65 years in age. Prakash happens to be the eldest while Kailash as youngest. As per the health history, it is estimated that the probability that Kailash survives another 5 years is 45\displaystyle 45\displaystyle \frac{4}{5}, Harish survives another 5 years is 35\displaystyle 35\displaystyle \frac{3}{5} and Prakash survives another 5 years is 12\displaystyle 12\displaystyle \frac{1}{2}. The probabilities that Kailash and Harish survive another 5 years is 0.46, Harish and Prakash survive another 5 years is 0.32 and Kailash and Prakash survive another 5 years is 0.48. The probability that all three sons survive another 5 years is 0.26. What shall be the probability that at least one of them survives another 5 years?

In a class, there are 15 boys and 10 girls. Three students are selected at random. The probability that 1 girl and 2 boys are selected is

Two cards are drawn from a pack of 52 cards. The probability that one is a spade and one is a heart; is

Ready to Master Probability?

Practice all 187 questions with instant feedback, earn XP, track your streaks, and ace your CA Foundation exam.

Start Practicing — It's Free